OpenStudy (anonymous):

Use the intercept method to graph the equation. 6x + 7y = 14

OpenStudy (caozeyuan):

ok so you want this in slope-intecept form which is y=mx+c where m is the slope and c is the intcept

OpenStudy (caozeyuan):

for obvious reasons. m=-6/7 and c=2 now you can graph it

OpenStudy (anonymous):

Okay when I go to solve I get (0,7/3) and (0,2) and when I enter it on my graph it doesn't work so I don't know what I am doing wrong

OpenStudy (caozeyuan):

No, don't solve

OpenStudy (caozeyuan):

In fact you cannot solve because funciton is different from eqations

OpenStudy (caozeyuan):

do you know how to graph a line if it is given in the form of y=mx+c

OpenStudy (anonymous):

No I don't understand and when I put in no solution I got that wrong so I am completely lost and confused

OpenStudy (caozeyuan):

calm down

OpenStudy (caozeyuan):

6x+7y=14 is the same as y=(-6/7)x+2 you understnad?

OpenStudy (anonymous):

no I don't see how you got that besides it looks like you combined the x and the y value

OpenStudy (caozeyuan):

\[6x+7y=14 \]\[7y=14-6x\]\[y=\frac{ 14-6x }{ 7 }\]\[y=2-\frac{ 6 }{ 7 }x\]

OpenStudy (caozeyuan):

better now?

OpenStudy (anonymous):

yes can I ask another question

OpenStudy (caozeyuan):

sure

OpenStudy (anonymous):

y-2x=-5/4

OpenStudy (anonymous):

I need three points

OpenStudy (caozeyuan):

well if you set y=0 you can get one

OpenStudy (caozeyuan):

x=0 gives you another one

OpenStudy (caozeyuan):

the third one can be y=1

OpenStudy (anonymous):

so when I use x=0 I get 5/4/1 which I don't know how to divide 1 by 5/4

OpenStudy (caozeyuan):

when x=0, y=-5/4

OpenStudy (caozeyuan):

so the point is (0, -5/4)

OpenStudy (anonymous):

so when x is 0 is the correct (0,0)

OpenStudy (caozeyuan):

but (0,0) is not on the line

OpenStudy (anonymous):

so what am I doing wrong

OpenStudy (caozeyuan):

y=2-(6/7)x specifies a relationship between x and y

OpenStudy (caozeyuan):

now we assign a value to x, in this case the value assigned is 0

OpenStudy (caozeyuan):

now from this x-value we can find the y-value becasue we have a relationship between x and y

OpenStudy (caozeyuan):

this relationship does not permit y to be 0 when x is 0

OpenStudy (anonymous):

so how do you solve 2-(6/7) and where did 6/7 come from

OpenStudy (anonymous):

If I am being dumb and annoying you you can stop helping me I just don't get math and it frustrates me

OpenStudy (caozeyuan):

I kind of want to say that

OpenStudy (caozeyuan):

I am going to sleep cause it is 3 in the morning and I got class tomorrow

OpenStudy (caozeyuan):

You should reread my posts

OpenStudy (anonymous):

Thanks

OpenStudy (caozeyuan):

the original function is 6x+7y=14

OpenStudy (caozeyuan):

then I use algebraic manipulation to rearrange the equation to the form y=mx+c, where m and c are constants

OpenStudy (caozeyuan):

then I plug given x-values or y-values ( 0, 1, etc) into the function

OpenStudy (caozeyuan):

now the function is a eqation and we can solve for he variable ( x or y)

OpenStudy (anonymous):

no fall break school all week boo hoo

OpenStudy (anonymous):

lets do y -4?